PHOTODISSOCIATION OF METHYL NITRITE FROM THE S2 STATE – ION IMAGES AND VECTOR CORRELATIONS

\underline{AIDAN J. KLOBUCHAR} and JEFFREY A. BARTZ, Department of Chemistry, Kalamazoo College, 1200 Academy Street, Kalamazoo, MI 49006.

Methyl nitrite, CH_{3}ONO, was photodissociated near 225 nm in a velocity-mapping time-of-flight mass spectrometer. The NO photoproduct was detected through state-resolved 1+1’ resonance-enhanced multiphoton ionization. The NO products are highly oriented in the laboratory frame, indicating a parallel transition and a fast dissociation.